# Leveraging the Bias-Variance Tradeoff in Quantum Chemistry for Accurate Negative Singlet-Triplet Gap Predictions: A Case for Double-Hybrid DFT

### v2

@@ 4 Results and Discussions
Figure 2: Error metrics for various methods using the aug-cc-pVDZ basis set to predict 12 theoretical best estimates of STGs for triangular molecules. MAD, MSE, and SDE represent mean absolute deviation, mean signed error, and standard deviation, respectively. Methods are ordered by increasing MAD, with a line included as a visual guide. CC3 and EOM-CCSD results are from Ref [14].
4.1 Overall trends across methods
Figure 2 presents an overview of the performance of various theoretical methods using the TBE of STGs from Ref. [14] as a reference. For consistency, all results are reported with the aug-cc-pVDZ basis set; results for other basis sets are provided in the Supplementary Information (SI). CC3 results are taken from Ref. [14]. Detailed error metrics for the wavefunction methods are provided in Table 1. For all methods listed in Table 1, increasing the basis set to aug-cc-pVTZ yields only minimal improvement. The reference (i.e., TBE) S1 energies were computed at the CC3/aug-cc-pVTZ level with a post-CC3 correction via EOM-CCSDT (using the 6-31+G(d) basis set), while the T1 energies were obtained at the EOM-CCSD/aug-cc-pVTZ level with a post-CCSD correction via CC3 using the aug-cc-pVDZ basis set [14]. Although the TBE approach treats S1 and T1 energies using different protocols, both are estimated at higher levels of theory currently available.
Among all methods, CC3/aug-cc-pVDZ shows the lowest values for MSE, MAD, and SDE ($-1$, 10, and 13 meV (milli-eV), respectively; see Table 1). When combined with the aug-cc-pVTZ basis set, CC3 essentially reproduces the TBE results. The next best performance is observed for L-CC2 with an MAD and SDE of 16 and 13 meV. L-ADC(2)/aug-cc-pVTZ results show an MAD and SDE of 22 and 13 meV, respectively, which are similar to the corresponding ADC(2) values of 20 and 16 meV.
Both the mean error and the spread of the error are significant for EOM-CCSD (Figure 2). Its MSE indicates that its predictions can serve as an upper bound for negative STGs; however, an SDE of approximately 80–90 meV diminishes confidence in predictions of gaps smaller than 0.1 eV, potentially leading to false negatives (i.e., systems with truly negative gaps might be predicted as having positive gaps). The large SDE suggests that the error is not systematic, hence, not all small, positive STGs predicted by EOM-CCSD may be reliably expected to be negative. In particular, the similarity-transformed local variant, DLPNO-STEOM-CCSD, exhibits a bias exceeding 0.25 eV with an SDE around 0.1 eV (see Figure 2), rendering it unsuitable for applications involving negative STG systems [68, 69].
Table 1: Error metrics for predicting twelve values of S1 & T1 energies, and STG of triangular molecules. For various methods, values are reported compared to the theoretical best estimates from Ref. [14]. MSE: mean signed error, MAD: mean absolute deviation, SDE: standard deviation of the error, minE: minimal error, and maxE: maximal error. In all cases, the basis set is aug-cc-pVDZ. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
Method Energy MSE MAD SDE minE maxE
L-CC2/aug-cc-pVDZa S1 0.047 0.047 0.016 0.020 0.071
T1 0.060 0.060 0.017 0.038 0.091
STG --0.013 0.016 0.013 --0.028 0.010
L-CC2/aug-cc-pVTZa S1 0.051 0.051 0.010 0.035 0.064
T1 0.065 0.065 0.016 0.032 0.090
STG --0.014 0.015 0.010 --0.028 0.005
L-ADC(2)/aug-cc-pVDZa S1 0.006 0.031 0.035 --0.056 0.056
T1 0.025 0.030 0.032 --0.020 0.078
STG --0.019 0.023 0.017 --0.038 0.023
L-ADC(2)/aug-cc-pVTZa S1 0.003 0.017 0.021 --0.037 0.039
T1 0.023 0.023 0.018 --0.002 0.052
STG --0.020 0.022 0.013 --0.036 0.009
ADC2/aug-cc-pVTZa S1 -0.000 0.021 0.026 -0.051 0.050
T1 0.015 0.022 0.022 -0.023 0.049
STG -0.016 0.020 0.016 -0.039 0.021
CC3/aug-cc-pVDZb S1 0.001 0.015 0.018 --0.030 0.026
T1 0.002 0.016 0.020 --0.038 0.037
STG --0.001 0.010 0.013 --0.035 0.015
  • a This work
  • b From Ref. [14]
Among the DFT methods, the dh-DFT approaches B2GP-PLYP and PBE-QIDH display very small SDEs (Figure 2), indicating that their deviations from the TBE values are primarily systematic. The spin-component-scaled (SCS) and opposite-spin-scaled (SOS) variants of these methods exhibit reduced MSE and MAD at the expense of larger SDEs. With few exceptions, the remaining DFAs show large errors, rendering them unreliable for first-principles identification of negative STG systems.
Figure 3: Bias-variance tradeoff in various methods as a scatterplot of MAD versus SDE for predicting 12 theoretical best estimates of STGs in triangular molecules. Wavefunction methods are marked with blue stars, and DFAs with red circles. For clarity, names are shown only for selected points. All results are based on the aug-cc-pVDZ basis set. CC3 and EOM-CCSD results are from Ref [14]
4.2 Bias-variance tradeoff across methods
@@
If the magnitude of negative STGs exceeds 0.2 eV, methods with mean errors $\approx 0.1$ eV may still allow for qualitative identification. However, for systems where the STG is below $\approx 0.1$ eV, similar prediction errors become too large to yield reliable qualitative conclusions. Figure 2 reveals interesting trends in the bias-variance tradeoff in various methods: while some methods exhibit a high MAD, they maintain a very low SDE. These cases are highlighted in Figure 3, which plots MAD versus SDE for various methods (excluding those with high MAD or SDE). Overall, CC3 and L-CC2 demonstrate superior performance with low MAD and SDE, whereas several DFT approaches—notably, B2GP-PLYP and PBE-QIDH—exhibit remarkably small variance. In particular, PBE-QIDH shows an SDE of less than 10 meV, indicating that its predictions are narrowly clustered around the TBE values but shifted by a constant systematic error. Although the MAD of PBE-QIDH exceeds 175 meV, systematic errors are generally easier to correct through linear regression than the non-systematic errors associated with high variance. For properties such as atomization or formation energies, DFT predictions are often corrected for systematic errors via dressed atom corrections, leading to lower errors for DFAs on the higher rungs of “Jacob’s ladder” [70, 56, 71].
Among the local correlated methods, L-CC2 exhibits MAD and SDE values close to those of CC3, while L-ADC2 shows a slightly larger SDE than L-CC2 (Figure 3). In contrast, the dh-DFT methods B2GP-PLYP and PBE-QIDH display SDEs that are smaller than L-CC2’s, but their MADs are larger by about an order of magnitude. To explore the potential of low-variance, high-bias methods, we applied a least-squares correction to the STG predictions of B2GP-PLYP, PBE-QIDH, L-ADC(2), and L-CC2 using TBE values as the reference; the results are collected in Table 2. As expected, the MSE becomes zero after linear correction. L-ADC(2) and L-CC2 show MADs of 0.023 and 0.016 eV with SDEs of 0.017 and 0.013 eV, respectively, with only slight improvement upon correction, indicating that their errors are primarily non-systematic. In contrast, dh-DFT methods B2GP-PLYP and PBE-QIDH, which initially exhibit MADs of 0.119 and 0.177 eV and SDEs of 0.011 and 0.009 eV respectively, experience a dramatic drop in MAD (to below 0.01 eV) after correction, while retaining their low SDEs. This implies that the errors in these dh-DFT methods are predominantly systematic. A similar analysis for the SCS and SOS variants of these methods, presented in Table 3, confirms that all fixed dh-DFT methods show reduced MAD and SDE values, indicating that their prediction errors are largely systematic when compared to TBE.
@@
Table 2: For various methods (aug-cc-pVDZ basis set) STGs of benchmark systems (Figure 1) are presented before and after bias-correction. Error metrics and the reference TBE values are also provided. MSE: mean signed error, MAE: mean absolute error, and SDE: standard deviation of the error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
# B2GP-PLYP PBE-QIDH L-ADC2 L-CC2 TBEe
@@
  • e From Ref. [14]
Table 3: Error metrics for predicting twelve values of S1 and T1 energetics of triangular molecules reported in Ref. [14]. Values are reported compared to the theoretical best estimates (TBE) from Ref. [14]. In all cases, the basis set is aug-cc-pVDZ, and the geometries are from Ref. [14]. MSE: mean signed error, MAE: mean absolute error, and SDE: standard deviation of the error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
@@
4.3 Internally referenced bias-variance correction for double-hybrid DFT
@@
While one can use L-CC2, CC3, or TBE values to correct PBE-QIDH predictions through linear scaling, we further investigated whether the low-variance DFT methods—B2GP-PLYP and PBE-QIDH—can be adjusted to minimize bias. The PBE-QIDH DFA is defined as follows [59]:
$E_{\rm xc}^{\rm PBE-QIDH}\left[\rho\right]$ $=$ $a_{x}E_{x}^{\rm HF}+\left(1-a_{x}\right)E_{x}^{\rm PBE}+a_{c}E_{c}^{\rm MP2}+\left(1-a_{c}\right)E_{c}^{\rm PBE}.$ (4.1)
@@
Here, the coefficients $a_{x}$ and $a_{c}$ control the fraction of exact exchange (via Hartree–Fock) and MP2-level correlation, respectively. In the standard formulation, PBE-QIDH ($a_{x}=0.69$ and $a_{c}=0.33$)[72] belongs to a family of dh-DFAs with $a_{c}=a_{x}^{3}$ delivering low errors for STG [73]. Similarly, B2GP-PLYP is based on the GGA-B88 exchange functional and a combination of LYP-GGA and VWN3-LDA correlation functionals:
$E_{\rm xc}^{\rm B2GP-PLYP}\left[\rho\right]$ $=$ $a_{x}E_{x}^{\rm HF}+\left(1-a_{x}\right)E_{x}^{\rm B88}+a_{c}E_{c}^{\rm MP2}+\left(1-a_{c}\right)\left(E_{c}^{\rm VWN3}+E_{c}^{\rm LYP}\right)$ (4.2)
with standard mixing coefficients $a_{x}=0.65$ and $a_{c}=0.36$ [74, 75, 76].
@@
Figure 4: Dependence of prediction errors of dh-DFT with exchange and correlation mixing parameters $a_{x}$ and $a_{c}$ (See Eq. 4.1 and Eq. 4.2). MSE, MAD and SDE (in meV) are shown for B2GP-PLYP (left panels) and PBE-QIDH (right panels) compared to TBE. Values of $a_{x}$ and $a_{c}$ were sampled in steps of 0.05 and the surfaces were smoothened using cubic functions. In all plots, the default mixing coefficients are marked by a red cross—(0.65, 0.36) for B2GP-PLYP and (0.69, 0.33) for PBE-QIDH—and coefficients that minimize MAD are marked by a blue star—(0.75, 0.55) for both B2GP-PLYP and PBE-QIDH.
@@
Prior studies have shown that the $a_{x}$-$a_{c}$ parameter space of B2GP-PLYP exhibits a continuous range of optimal parameters depending on the target property [58, 23]. For instance, Tarnopolsky et al. demonstrated variations in the SDE for thermochemistry and reaction barriers as a function of $a_{x}$ and $a_{c}$, and proposed the B2T-PLYP and B2K-PLYP df-DFAs with ($a_{x}=0.60$, $a_{c}=0.31$) and ($a_{x}=0.72$, $a_{c}=0.40$), respectively [22]. However, while these studies focused on a single error metric across different properties—emphasizing variance because bias in thermochemistry is corrected through quasi-atomic corrections[70, 56, 71]—the bias-variance trade-off for a single property such as the STG has not been thoroughly examined.
@@
Using B2GP-PLYP and PBE-QIDH DFAs, we scanned $a_{x}$ and $a_{c}$, and computed error metrics for all 12 benchmark STGs using TBE as the reference. As shown in Figure 4, both dh-DFAs exhibit low SDEs for the default mixing parameters, and a continuous range of ($a_{x}$, $a_{c}$) combinations yielding zero MSE is observed (Figure 4, top panels). Some calculations failed to converge when $a_{x}$ and $a_{c}\approx 0$; these are not shown in Figure 4. However, the minimum MAD is achieved at $a_{x}=0.75$ and $a_{c}=0.55$, with both DFAs delivering MADs of approximately 25 meV compared to TBE. We denote these optimized parameterizations as B2GP-PLYP (75,55) and PBE-QIDH (75,55). To assess the variation of optimal $a_{x}$ and $a_{c}$ across error metrics for S1 and T1 energies separately, we performed a comparative analysis (Table S1) using PBE-QIDH. Notably, increased accuracy in S1 energies is observed when both mixing parameters approach 1 (with $a_{x}=0.90$ and $a_{c}=0.85$ yielding a MAD of 68 meV), while the lowest MAD for T1 (22 meV) is obtained for $a_{x}=0.5$ and $a_{c}=0.15$.
To apply the scaling protocol discussed in Section 3, the default B2GP-PLYP and PBE-QIDH parameterizations serve as the low-variance estimators ($y_{1}$ in Eq.(3.1)), while the (75,55) variants, which exhibit the lowest MAD, serve as the reference low-bias estimators ($y_{2}$ in Eq.(3.1)). We then determine the slope and intercept via linear regression to correct the bias, with the corrected predictions denoted as B2GP-PLYP and PBE-QIDH. Table 4 presents the STG values for the 12 benchmark systems predicted by the bias-corrected dh-DFAs. Importantly, the scaling procedure does not require any additional high-level reference data to identify the low-bias or low-variance domain in the parameter space of dh-DFAs. The bias-corrected predictions, which simultaneously reflect the small SDE of the base dh-DFAs and the low MAD of the internal references, achieve MAD||||SDE values of 10||||12 and 9||||11 meV for B2GP-PLYP and PBE-QIDH, respectively, with respect to TBE.
@@
Figure 5: Comparison of 12 STGs of benchmark systems across methods: a) Scatterplot of values predicted by various methods (with the aug-cc-pVDZ basis set) versus theoretical best estimates (TBE). TBE and CC3 results are from Ref. [14]. b) Distribution of errors in predicted STGs relative to TBE shown as smooth densities.
@@
The overall correlation between the predicted values and TBE is illustrated in Figure 5. Due to the large bias in the original dh-DFT predictions, the Pearson correlation coefficients are strongly negative ($\rho=-0.61$ for B2GP-PLYP and $\rho=-2.45$ for PBE-QIDH; see Figure 5a). Upon applying the internally referenced scaling, the correlation improves markedly to $\rho=0.98$ for both DFAs, with all 12 values aligning with TBE more closely than those from L-ADC(2), L-CC2, or even CC3. The overall performance of the scaling procedure stems from the exceptionally small variance of the dh-DFAs and the low MAD achieved with the (75,55) parameterizations.
4.4 Application to heptazine derivatives

### v3

@@ IV.1 Performance of L-CC2 and L-ADC(2)
Figure 2: Error metrics for various methods using the aug-cc-pVDZ basis set to predict 12 theoretical best estimates of STGs for triangular molecules. MAD, MSE, and SDE represent mean absolute deviation, mean signed error, and standard deviation, respectively. Methods are ordered by increasing MAD, with a line included as a visual guide. CC3 and EOM-CCSD results are from Ref. Loos, Lipparini, and Jacquemin, 2023.
IV.2 Overall trends across methods
Figure 2 presents an overview of the performance of various theoretical methods, using as a reference the TBEs of STGs collected from Ref. Loos, Lipparini, and Jacquemin, 2023. In this figure, MSE, MAD, and SDE are plotted for all methods, sorted by increasing values of MAD. In this analysis, we also included CC3/aug-cc-pVDZ results from Ref. Loos, Lipparini, and Jacquemin, 2023. For consistency, Figure 2 presents results based on the aug-cc-pVDZ basis set, while similar figures for other basis sets are provided in the Supplementary Information (SI).
Among all methods, CC3/aug-cc-pVDZ shows the lowest values for MSE, MAD, and SDE ($-1$, 10, and 13 meV (milli-eV), respectively). As these errors are an order of magnitude smaller than the typical magnitudes of reported negative STGs, CC3/aug-cc-pVDZ can be considered as a computationally affordable reference method. The next best performance is observed for L-CC2, with MSE/MAD/SDE values of --13/16/13 meV, respectively.
The MSEs of CC2, L-CC2, ADC(2), L-ADC(2), and CC3 are negative, indicating that these methods will prefer slightly more negative STGs compared to the reference TBE values. Such a systematic trend will result in false-positive predictions of DFIST systems, especially when the magnitudes of the STGs are comparable or smaller than the mean errors of the methods. However, if the actual STGs of the DFIST candidate molecules are lower than the MAD of the corresponding method (such as $\lesssim-15$ meV for L-CC2/aug-cc-pVDZ), then predictions can be considered as true-positives.
EOM-CCSD method is another popular wavefunction method applied to identify DFIST candidate molecules. The local correlated version of the method DLPNO-STEOM-CCSD has been very successful in yielding CCSD quality results with attractive speedups. In Table 3, we present the excited state energies and STGs of the benchmark set obtained with the EOM-CCSD and DLPNO-STEOM-CCSD methods with the aug-cc-pVDZ basis set.
Table 3: EOM-CCSD energies and DLPNO-STEOM-CCSD energies with aug-cc-pVDZ of the S1 and T1 states with respect to the S0 ground state along with the singlet-triplet gap, S1-T1, of 12 triangular benchmark systems. All values are in eV and # indicates compound number as in Figure 1.
# EOM-CCSDa DLPNO-STEOM-CCSDb
S1 T1 S1-T1 S1 T1 S1-T1
1 $2.901$ $3.049$ $-0.148$ $2.243$ $2.930$ $-0.687$
2 $1.092$ $1.165$ $-0.073$ $0.662$ $1.093$ $-0.431$
3 $1.703$ $1.732$ $-0.029$ $1.297$ $1.649$ $-0.352$
4 $2.342$ $2.382$ $-0.040$ $1.790$ $2.275$ $-0.485$
5 $2.290$ $2.312$ $-0.022$ $1.800$ $2.181$ $-0.381$
6 $0.932$ $0.944$ $-0.012$ $0.519$ $0.888$ $-0.369$
7 $0.775$ $0.761$ $+0.014$ $0.393$ $0.685$ $-0.292$
8 $0.619$ $0.595$ $+0.024$ $0.246$ $0.527$ $-0.281$
9 $1.388$ $1.511$ $-0.123$ $0.740$ $1.362$ $-0.622$
10 $1.666$ $1.907$ $-0.241$ $0.962$ $1.747$ $-0.785$
11 $0.864$ $0.989$ $-0.125$ $0.312$ $0.850$ $-0.538$
12 $1.415$ $1.172$ $+0.243$ $0.968$ $1.020$ $-0.052$
MSE +0.131 +0.054 +0.077 $-0.374$ $-0.055$ $-0.319$
MAD +0.131 +0.054 +0.077 +0.374 $+0.055$ +0.319
SDE +0.042 +0.031 +0.033 +0.102 +0.042 +0.100
  • a From Ref. Loos, Lipparini, and Jacquemin, 2023.
  • b This work
Among the wavefunction methods considered in this study, EOM-CCSD is the only method delivering a positive MSE (see Table 3 and also Figure 2). This trend suggests that EOM-CCSD predictions are prone to false-negative predictions in identifying systems with negative STGs, where an actual negative STG can be incorrectly forecasted as positive. There is a slight advantage in using EOM-CCSD as its predictions of STGs can serve as upper bounds for negative STGs, while the actual values can be considered smaller. However, the SDE of EOM-CCSD is $>30$ meV, suggesting that the positive MSE of this method is not systematic. Hence, not all small, positive STGs predicted by EOM-CCSD may be reliably expected to be negative.
DLPNO-STEOM-CCSD has demonstrated to predict the STGs of multi-resonance TADF compounds in good agreement with EOM-CCSD and ADC(3) Izu, Matxain, and Casanova (2024), as well as vertical excitation energies of organic radicals aligning with EOM-CCSD, CC3, and CCSDT Casanova-Páez and Neese (2024). On the otherhand, for the twelve benchmark systems explored here, DLPNO-STEOM-CCSD, exhibits a bias (MAD) exceeding 0.3 eV with an SDE of 0.1 eV (see Table 3 and Figure 2), rendering it unsuitable for applications involving negative STG systems concurring with previous studies Chanda and Sen (2024); Chanda, Saha, and Sen (2025). To understand the origin of this error, let us consider System 1 with the TBE values of the excitation energies of S1 and T1 states reported as 2.717 eV and 2.936 eV, resulting in an STG of $-0.219$ eV (see Table 1). At the EOM-CCSD/aug-cc-pVDZ level, these values were predicted as 2.901 eV (S1), 3.049 eV (T1), and $-0.148$ eV (STG), see Table 3. However, with DLPNO-STEOM-CCSD/aug-cc-pVDZ, the same values were predicted as 2.243 eV (S1), 2.930 eV (T1), and $-0.687$ eV (STG), see Table 3. This comparison suggests that the T1 energies of DLPNO-STEOM-CCSD agree with those of canonical EOM-CCSD, while the former method significantly lowers the S1 energy, resulting in more negative STGs.
Among the DFT methods, only dh-DFAs can predict negative STGs within the LR-TDDFT formalism Sancho-Garcia et al. (2022); Casanova-Páez and Goerigk (2021a). However, it must be noted that studies have also shown that negative STGs can be predicted using hGGA methods when calculating the S1 and T1 states using the $\Delta$-SCF formalism based on the maximal overlap method Kunze et al. (2024); Chanda, Saha, and Sen (2025). In Figure 2, one finds that the dh-DFAs SOS-PBE-QIDH and SOS-$\omega$PBEPP86 have MADs very similar to those of L-CC2, ADC(2), and L-ADC(2). Especially for these two methods, the MSEs are nearly zero, indicating that they intrinsically incorporate bias correction in their parameterization. Further, the dh-DFT approaches B2GP-PLYP and PBE-QIDH display very small SDEs (Figure 2), indicating that their deviations from the TBE values are primarily systematic. This is a key result that will be explored further in Section IV.3. The SCS and SOS variants of these methods exhibit smaller MSEs and MADs at the expense of larger SDEs. With few exceptions, the remaining DFAs exhibit significantly larger errors, rendering them unreliable for the first-principles identification of negative STG systems when used within the LR-TDDFT formalism.
IV.3 Bias-variance tradeoff across methods
@@
While interesting trends in the bias-variance tradeoff of various methods are noticeable in Figure 2, the trend is more clearly revealed in Figure 3, which displays a scatterplot of MADs and SDEs for various methods (excluding those with high MAD or SDE). Overall, CC3 and L-CC2 demonstrate superior performance, characterized by low MAD and SDE. The dh-DFT methods, B2GP-PLYP and PBE-QIDH, exhibit high MADs while maintaining very low SDEs. Such methods are suitable for bias correction, denoted as $\hat{y}_{1}$ in Section III.2. In particular, PBE-QIDH shows an SDE of less than 10 meV, indicating that its predictions are narrowly clustered around the TBE values but shifted by a constant systematic error. Although the MAD of PBE-QIDH exceeds 175 meV, systematic errors are generally easier to correct through linear regression than the non-systematic errors associated with high variance.
Figure 3: Bias-variance tradeoff in various methods as a scatterplot of MAD versus SDE for predicting 12 theoretical best estimates of STGs in triangular molecules. Wavefunction methods are marked with blue stars, and DFAs with red circles. For clarity, names are shown only for selected points. All results are based on the aug-cc-pVDZ basis set. CC3 and EOM-CCSD results are from Ref. Loos, Lipparini, and Jacquemin, 2023.
@@
Table 4: For various methods (aug-cc-pVDZ basis set) STGs of benchmark systems (Figure 1) are presented before and after bias-correction. Error metrics and the reference TBE values are also provided. MSE: mean signed error, MAD: mean absolute deviation, and SDE: standard deviation of the error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
# B2GP-PLYP PBE-QIDH L-ADC(2) L-CC2 TBEe
@@
  • e From Ref. Loos, Lipparini, and Jacquemin, 2023
Among the local correlated methods, L-CC2 exhibits MAD and SDE values close to those of CC3, while L-ADC(2) shows a slightly larger SDE than L-CC2 (Figure 3). In contrast, the dh-DFT methods B2GP-PLYP and PBE-QIDH display SDEs that are smaller than L-CC2’s, but their MADs are larger by about an order of magnitude. To explore the potential of such low-variance, high-bias methods, we applied a least-squares correction to the STG predictions of B2GP-PLYP, PBE-QIDH, L-ADC(2), and L-CC2 using TBE values as the reference. These results are collected in Table 4. For all methods, the MSEs become zero after linear correction as demonstrated in Section III.3.
L-ADC(2) and L-CC2 exhibit MAD||||SDEs of $0.023|0.017$ and $0.016|0.013$ eV, respectively, which slightly improve to $0.014|0.016$ and $0.010|0.012$ eV upon linear corrections referenced to the TBE values. The dh-DFT methods B2GP-PLYP and PBE-QIDH, which initially exhibit MAD/SDEs of $0.119|0.011$ and $0.177|0.009$ eV, retain their SDE upon linear correction, but their MADs drop to less than 0.01 eV. This implies that the errors in the STGs predicted by these dh-DFT methods are predominantly systematic in nature.
Table 5: Error metrics for predicting twelve values of S1 and T1 energetics of triangular molecules reported in Ref. Loos, Lipparini, and Jacquemin, 2023. Values are reported compared to the theoretical best estimates (TBE) from Ref. Loos, Lipparini, and Jacquemin, 2023. In all cases, the basis set is aug-cc-pVDZ, and the geometries are from Ref. Loos, Lipparini, and Jacquemin, 2023. MSE: mean signed error, MAD: mean absolute deviation, and SDE: standard deviation of the error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
@@
IV.4 Internally referenced bias-variance correction for double-hybrid DFT
@@
While one can use L-CC2, CC3, or TBE values to correct PBE-QIDH predictions through linear scaling, we further investigated whether the low-variance DFT methods, B2GP-PLYP and PBE-QIDH, can be corrected to minimize their systematic errors (or intrinsic bias) discussed above. The PBE-QIDH DFA is defined as follows Brémond et al. (2014)
$E_{\rm xc}^{\rm PBE-QIDH}\left[\rho\right]$ $=$ $a_{x}E_{x}^{\rm HF}+\left(1-a_{x}\right)E_{x}^{\rm PBE}+a_{c}E_{c}^{\rm MP2}$ (11)
$+\left(1-a_{c}\right)E_{c}^{\rm PBE}.$
@@
Here, the coefficients $a_{x}$ and $a_{c}$ control the fraction of exact exchange (via Hartree–Fock) and MP2-level correlation, respectively. In the standard formulation, PBE-QIDH ($a_{x}=0.69$ and $a_{c}=0.33$)Sandoval-Salinas et al. (2023) belongs to a family of dh-DFAs with $a_{c}=a_{x}^{3}$ delivering low errors for STG Derradji et al. (2024).
Similarly, B2GP-PLYP is based on the GGA-B88 exchange functional combined with LYP-GGA and VWN3-LDA correlation functionals
$E_{\rm xc}^{\rm B2GP-PLYP}\left[\rho\right]$ $=$ $a_{x}E_{x}^{\rm HF}+\left(1-a_{x}\right)E_{x}^{\rm B88}+a_{c}E_{c}^{\rm MP2}$ (12)
$+\left(1-a_{c}\right)\left(E_{c}^{\rm VWN3}+E_{c}^{\rm LYP}\right)$
with standard mixing coefficients $a_{x}=0.65$ and $a_{c}=0.36$ Kozuch, Gruzman, and Martin (2010); Goerigk, Moellmann, and Grimme (2009); Pantazis (2019).
@@
Figure 4: Dependence of prediction errors of dh-DFT with exchange and correlation mixing parameters $a_{x}$ and $a_{c}$ (See Eq. 11 and Eq. 12). MSE, MAD and SDE (in meV) are shown for B2GP-PLYP (left panels) and PBE-QIDH (right panels) compared to TBE. Values of $a_{x}$ and $a_{c}$ were sampled in steps of 0.05 and the surfaces were smoothened using cubic functions. In all plots, the default mixing coefficients are marked by a red cross—(0.65, 0.36) for B2GP-PLYP and (0.69, 0.33) for PBE-QIDH—and coefficients that minimize MAD are marked by a blue star—(0.75, 0.55) for both B2GP-PLYP and PBE-QIDH.
@@
Prior studies have shown that the $a_{x}$-$a_{c}$ parameter space of B2GP-PLYP exhibits a continuous range of optimal parameters depending on the target property Karton et al. (2008); Goerigk and Grimme (2011). For instance, Tarnopolsky et al. demonstrated variations in the SDE for thermochemistry and reaction barriers as a function of $a_{x}$ and $a_{c}$, and proposed the B2T-PLYP and B2K-PLYP df-DFAs with ($a_{x}=0.60$, $a_{c}=0.31$) and ($a_{x}=0.72$, $a_{c}=0.40$), respectively Tarnopolsky et al. (2008). However, while these studies focused on a single error metric across different properties, mainly emphasizing variance because bias in thermochemistry is corrected through quasi-atomic correctionsWinget and Clark (2004); Grimme (2006); Das, Chakraborty, and Ramakrishnan (2021), the bias-variance trade-off for a single property, such as the STG, has not been thoroughly examined.
@@
Using B2GP-PLYP and PBE-QIDH DFAs, we scanned the $a_{x}$-$a_{c}$ space and computed error metrics for all 12 benchmark STGs using TBE as the reference. As shown in Figure 4, both dh-DFAs exhibit low SDEs for the default mixing parameters, and a continuous range of ($a_{x}$, $a_{c}$) combinations yielding zero MSE (Figure 4, top-most panels). Some calculations failed to converge when $a_{x}$ and $a_{c}\approx 0$; these are not shown in Figure 4. However, the minimum MAD is achieved at $a_{x}=0.75$ and $a_{c}=0.55$, with both DFAs delivering MADs of approximately 25 meV compared to TBE. We denote these optimized parameterizations as B2GP-PLYP (75,55) and PBE-QIDH (75,55). To assess the variation of optimal $a_{x}$ and $a_{c}$ across error metrics for S1 and T1 energies separately, we performed a comparative analysis (Table S1 in the SI) using PBE-QIDH. Notably, increased accuracy in S1 energies is observed when both mixing parameters approach 1 (with $a_{x}=0.90$ and $a_{c}=0.85$ yielding a MAD of 68 meV), while the lowest MAD for T1 (22 meV) is obtained for $a_{x}=0.5$ and $a_{c}=0.15$.
@@
Figure 5: Comparison of 12 STGs of benchmark systems across methods: a) Scatterplot of values predicted by various methods (with the aug-cc-pVDZ basis set) versus theoretical best estimates (TBE). TBE and CC3 results are from Ref. Loos, Lipparini, and Jacquemin, 2023. b) Distribution of errors in predicted STGs relative to TBE shown as smooth densities.
To apply the scaling protocol discussed in Section III, the default B2GP-PLYP and PBE-QIDH parameterizations serve as the low-variance estimators ($y_{1}$ in Eq. 1), while the (75,55) variants, which exhibit the lowest MAD, serve as the reference low-bias estimators ($y_{2}$ in Eq. 1). We then determine the slope and intercept via linear regression to correct the bias, and denote the corrected predictions as B2GP-PLYP and PBE-QIDH.
Table 6 presents the STG values for the 12 benchmark systems predicted by the bias-corrected dh-DFAs. Importantly, the scaling procedure does not require any additional high-level reference data to identify the low-bias or low-variance domain in the parameter space of dh-DFAs. The bias-corrected predictions, which simultaneously reflect the small SDE of the base dh-DFAs and the low MAD of the internal references, achieve MAD/SDE values of 10/12 and 9/11 meV for B2GP-PLYP and PBE-QIDH, respectively, with respect to TBE.
@@
The overall correlation between the predicted values and TBE is illustrated in Figure 5. Due to the large bias in the original dh-DFT predictions as revealed in Figure 5a, their Pearson correlation coefficients are strongly negative ($\rho=-0.61$ for B2GP-PLYP and $\rho=-2.45$ for PBE-QIDH). Upon applying the internally referenced scaling, the correlation improves markedly to $\rho=0.98$ for both DFAs, with all 12 values aligning more closely with the TBE values than those from L-ADC(2), L-CC2, or even CC3. The overall performance of the scaling procedure stems from the exceptionally small variance of the dh-DFAs with the default parametrization and the low MAD achieved by the (75,55) parameterizations.
IV.5 Application to heptazine derivatives